Solve the linear system: 6x+6y= -24 , 2x+3y=-13
Answer to choose from
The solution is : (-1,5)
The solution is : (-2,5)
The solution is : (-3,5)
The solution is : (-4,5)

Answer by mananth(16949)

6x+6y=-24 .............1
2x+3y= -13 .............2

multiply (1)by 1
Multiply (2) by -3
6x+6y=-24
-6x-9y=39

Add the two equations
-3y=15

/-3

y=-5
Plug value of y in
6x+6y=-24
6x-30=-24
6x=6
/6
x=1